# Test-Time Modification: Inverse Domain Transformation for Robust Perception

### v2

@@ 4.2 Object Detection
A performance comparison for the domain-generalized object detection task on Cityscapes $\rightarrow$ BDD100 K Night-Det is shown in Table 4. We report the mean Average Precision (mAP) metric at $0.5$ threshold. We find that TTM yields notable gains by doubling and even tripling the mAP score for Faster-RCNN and Mask R-CNN, respectively. Qualitative results are shown in Fig. 7. As is visually evident from the figure, detecting objects becomes much easier after TTM, as challenging nighttime conditions are turned into easier daytime ones.
Table 5: Image classification performance comparison on ImageNet-R dataset. Top-1 accuracies are reported in %.
Method ImageNet-R (%) $\Delta\uparrow$ (%)
ResNet-50 36.1
   + ImageNet-21K Pretraining 37.2 +3.0%
   + CBAM (Self-Attention) 36.8 +1.9%
   + $\ell_{\infty}$ Adversarial Training 31.4 -13.0%
   + Speckle Noise 37.9 +5.0%
   + Style Transfer Augmentation 41.5 +15.0%
   + AugMix 41.1 +13.8%
   + DeepAugment 42.2 +16.9%
   + DeepAugment + AugMix 46.8 +29.6%
ResNet-50 + TTM (Ours) 60.8 +68.5%
ResNet-152 41.3
ResNet-152 + TTM (Ours) 63.5 +53.8%
@@ 4.3 Image Classification
For image classification task, apart from the base model (w/o TTM), we also compare against other models introduced in [hendrycks2021many] that are trained on ImageNet-1K while using additional strategies to reduce the domain gap like pre-training on ImageNet-21K (10x the size of ImageNet-1K), different augmentation techniques, using self-attention for better spatial feature learning, and adversarial training. However, we only take raw ResNet-50 and ResNet-152 models pre-trained on ImageNet-1K as candidates for applying TTM without using any of the additional domain reduction techniques. Here, due to the large validation set size, for image generation, we use a 4-bit quantized variant of QIE-2509 with 8-step Lightning LoRA weights for faster inference. The improvement should be even higher if we use the non-quantized complete QIE-2509 model.
Quantitative results with our TTM method are shown in Table 5. As can be seen, we are able to get the highest top-1 accuracy with ResNet-50 using TTM on ImageNet-R without employing any additional training strategies. Moreover, even with the larger ResNet-152 model, we are able to significantly improve the top-1 accuracy. Additionally, following a similar trend to domain generalized semantic segmentation, applying TTM with ResNet-50 even surpasses the performance of a larger ResNet-152 model (from 41.3 to 60.8). Qualitative results are shown in Fig. 8. As visible from the figure, the transformation of different renditions to real-life-like objects reduces the domain gap w.r.t ImageNet thereby helping to achieve better image classification performance for a fixed discriminative model.
@@ 5 Conclusion
In this paper, we propose an efficient approach to improve the perception robustness during inference time called Test-Time Modification (TTM). During TTM, we perform an inverse domain transformation of the target domain images back to the source domain by implicitly distilling the world knowledge of state-of-the-art Image-to-Image generation models. To demonstrate the effectiveness of our approach, we experimentally focus on the domain generalization setting for semantic segmentation, object detection, and image classification tasks, utilizing challenging target data distributions across various weather and daytime conditions, as well as rendering shifts. The significant improvements in our quantitative and qualitative results across multiple downstream tasks, datasets, and generation models demonstrate that test-time modification using large, information-rich Image-to-Image generation models can significantly help in creating more robust perception pipelines for challenging environments.
6 Acknowledgment
The research is funded by the German Federal Ministry for Economic Affairs and Energy within the project “NXT GEN AI METHODS – Generative Methoden für Perzeption, Prädiktion und Planung”. We thank the consortium and also acknowledge the Gauss Centre for Supercomputing for providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC).

### v4

@@ 4.2 Object Detection
A performance comparison for the domain-generalized object detection task on Cityscapes $\rightarrow$ BDD100 K Night-Det is shown in Tab.˜5. We report the mean Average Precision (mAP) metric at $0.5$ threshold. We find that TTM yields notable gains by doubling and even tripling the mAP score for Faster-RCNN and Mask R-CNN, respectively. Qualitative results are shown in Fig.˜7. As is visually evident from the figure, detecting objects becomes much easier after TTM, as challenging nighttime conditions are turned into easier daytime ones.
Figure 8: Segmentation results with TTM on BDD100K-Night and ACDC using a Cityscapes pretrained Mask2Former: comparing large and small I2I models.
I2I Generation Model $^{\text{BDDNight}} mIoU$$^{\text{ACDC}}$ Base Model (w/o TTM) 40.6 60.5 Large Flux.1 Kontext Dev [8 steps] 40.1 57.0 QIE-2509 47.1 66.4 Flux.1 Kontext Max 49.1 64.3 Small Flux.2 Klein Base 9B [4 steps] 50.9 67.7 Flux.2 Klein Base 4B [4 steps] 50.2 65.1 Flux.2 Klein 9B 49.9 67.4 Flux.2 Klein 4B 50.9 64.0
Figure 9: Image classification performance comparison on ImageNet-R dataset. Top-1 accuracies are reported.
Method ImageNet-R (%) $\Delta\uparrow$ ResNet-50 36.1 + ImageNet-21K Pretraining 37.2 +1.1 + CBAM (Self-Attention) 36.8 +0.7 + $\ell_{\infty}$ Adversarial Training 31.4 -4.7 + Speckle Noise 37.9 +1.8 + Style Transfer Augmentation 41.5 +5.4 + AugMix 41.1 +5 + DeepAugment 42.2 +6.1 + DeepAugment + AugMix 46.8 +10.7 ResNet-50 + TTM (Ours) 60.8 +24.7 ResNet-152 41.3 ResNet-152 + TTM (Ours) 63.5 +22.2
@@ 4.3 Image Classification
Beyond our main semantic segmentation and object detection tasks, we additionally analyze TTM on an image classification task to assess its behavior in a substantially different setting. For image classification task, apart from the base model (w/o TTM), we also compare against other models introduced in [12] that are trained on ImageNet-1K while using additional strategies to reduce the domain gap like pre-training on ImageNet-21K (10x the size of ImageNet-1K), different augmentation techniques, using self-attention for better spatial feature learning, and adversarial training. However, we only take raw ResNet-50 and ResNet-152 models pre-trained on ImageNet-1K as candidates for applying TTM without using any of the additional domain reduction techniques. Here, due to the large validation set size, for image generation, we use a 4-bit quantized variant of QIE-2509 with 8-step Lightning LoRA weights for faster inference. The improvement should be even greater if we use the non-quantized complete QIE-2509 model.
Quantitative results with our TTM method are shown in Fig.˜9. As can be seen, we are able to get the highest top-1 accuracy with ResNet-50 using TTM on ImageNet-R without employing any additional training strategies. Moreover, even with the larger ResNet-152 model, we are able to significantly improve the top-1 accuracy. Additionally, following a similar trend to domain generalized semantic segmentation, applying TTM with ResNet-50 even surpasses the performance of a larger ResNet-152 model (from 41.3 to 60.8). Qualitative results are shown in Fig.˜10. As visible from the figure, the transformation of different renditions to real-life-like objects reduces the domain gap w.r.t ImageNet thereby helping to achieve better image classification performance for a fixed discriminative model.
4.3.1 Inference Time and Deployment.
TTM introduces a modular image transformation step that operates independently of the downstream perception model. We analyze the runtime characteristics of the TTM pipeline, focusing on two components: prompt generation via MLLM metaprompting and the inference speed of the I2I generation model.
MLLM metaprompting: We employ a single task-level prompt shared across all images. As a result, the MLLM is invoked only once to generate the source-domain prompt, and this step is detached from the per-image inference pipeline.
I2I generation: The main computational step of TTM is the I2I transformation applied to each test image. Recent advances in efficient generative architectures substantially reduce runtime and memory requirements while maintaining high generation quality. Figure˜9 compares the post-TTM segmentation performance of lightweight and large I2I models on the BDDNight and ACDC datasets. The results show that reducing model complexity does not compromise downstream performance.
[…]
As illustrated in Fig.˜11, modern accelerators further accelerate this process through higher throughput and low-precision execution. For example, Flux.2 Klein 4B generates an image in 2s on an A100 GPU, 0.9s on an H100, 0.8s on an H200, and 0.4s on a B200 GPU. These results demonstrate that TTM can operate at practical inference speeds on modern hardware. Additional improvements through batching, quantization, and emerging inference accelerators (e.g., Groq LPU [1], Cerebras WSE [19]) are expected to further reduce latency. Additional runtime analysis of the I2I generation models within the TTM framework is included in the appendix.
Figure 11: Average generation time per image for different I2I generation models across different GPUs in seconds per image. VRAM occupancy for each model is also depicted on the right. * bar refers to ASIC chip [1, 19] inference time estimates using public benchmarks. Note that we run Flux.2 Klein Base (4B/9B) for just 4 steps, as this already produces satisfactory images for TTM. We already reach near real-time inference with smaller models.
@@ 5 Conclusion
In this paper, we propose an efficient approach to improve the perception robustness during inference time called Test-Time Modification (TTM). During TTM, we perform an inverse domain transformation of the target domain images back to the source domain by implicitly distilling the world knowledge of state-of-the-art Image-to-Image generation models. To demonstrate the effectiveness of our approach, we experimentally focus on the domain generalization setting for semantic segmentation, object detection, and image classification tasks, utilizing challenging target data distributions across various weather and daytime conditions, as well as rendering shifts. The significant improvements in our quantitative and qualitative results across multiple downstream tasks, datasets, and generation models demonstrate that test-time modification using large, information-rich Image-to-Image generation models can significantly help in creating more robust perception pipelines for challenging environments. Finally, our deployment analysis shows that these gains are achievable with realistic latencies today, and that continued advances in efficient architectures and accelerator hardware make real-time deployment increasingly feasible.
Acknowledgements
The research is funded by the German Federal Ministry for Economic Affairs and Energy within the project “NXT GEN AI METHODS – Generative Methoden für Perzeption, Prädiktion und Planung". We thank the consortium and also acknowledge the Gauss Centre for Supercomputing for providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC).